solve the rational inequality and graph the solution set on a real number line.
x+12
----- < 3
x-9

Answer by stanbon(75887)

x+12
----- < 3
x-9
-------------
Solve the equality:
x+12 = 3(x-9)
x+12 = 3x-27
2x = 39
x = 19.5
x cannot be 9; x cannot be -12; x cannot be 19.5
----------------------
Draw a number line and mark -12 and 9 and 19.5 on it.
-------------
Check a test point in each resulting interval to find the solution set.
--------------
Let x = -13, you get (-13+12)/(-13-9)<3; true so solutions in (-inf,-12)
Let x = 0, you get (0+12)/(0-9) <3; true so solutions in (-12,9)
Let x = 10, you get (10+12)/(10-9)<3 ; false so no solutions in (9,19.5)
Let x = 20, you get (20+12)/(20-9)<3 ; true so solutions in (19.5,+inf)
------------------
Final Answer: (-inf,-12)U(-12,9)U(19.5,+inf)
======================
Cheers,
Stan H.